# Surface groups among cubulated hyperbolic and one-relator groups

### v1

@@ 2.3 Connected sums of simplicial complexes
2.4 Cut-set lemmas in connnected sums
@@
Again, let $C=A\#_{v}B$ for notational convenience. The hypothesis that $\sigma$ is a cut simplex of $C$ means that $C_{\sigma}$ is disconnected. The proof now divides into cases, according to whether $\sigma$ is in $A$ but not in $B$, or $\sigma$ is contained in the canonical copy of the link of $v$.
@@
so Lemma 2.22 implies that $v$ is a cut vertex in $B$, as in item (i).
@@ 2.5 Cuts and the generalised Whitehead’s lemma
Consider a finite CAT(0) cube complex equipped with an immersion $Y\to X$. If $\mathrm{Wh}_{X}(Y)$ has a finite cut set then $Y$ is said to be a cut. More precisely, if $\mathrm{Wh}_{X}(Y)$ is $(k-1)$-connected but not $k$-connected – i.e. if the smallest cut set for $\mathrm{Wh}_{X}(Y)$ has cardinality $k$ – then $Y$ is said to be a $k$-cut. If $\{H_{1},\ldots,H_{k}\}$ is a minimal cut set for a $k$-cut $Y$ then we say that $Y$ is a cut between $H_{1},\ldots,H_{k}$.
@@
Consider a finite CAT(0) cube complex equipped with a locally isometric immersion $Y\to X$, and let $Y=Y_{1}\cup_{H}Y_{2}$ for some hyperplane $H$ of $Y$. If $Y$ is a 0-cut then either one of $Y_{1}$ or $Y_{2}$ is a 0-cut, or both $Y_{1}$ and $Y_{2}$ are 1-cuts.
@@
In the base case, $Y$ is a single vertex $y$, so $\mathrm{Wh}_{X}(Y)\cong\mathrm{Lk}_{X}(y)$ by Example 2.6 and there is nothing to prove. If $Y$ has more than one vertex, we may choose a (necessarily separating) hyperplane $H$ of $Y$ and write $Y=Y_{1}\cup_{H}Y_{2}$. By the inductive hypothesis, both $Y_{1}$ and $Y_{2}$ are connected without cut simplices, and therefore so is $Y$ by Lemma 2.27.
@@
Theorem E provides a convenient certificate that $\pi_{1}(X)$ does not split freely. However, the necessary condition that it provides is not sufficient: there are non-positively curved square complexes in which the link of a vertex contains a cut vertex, and yet the universal cover is one-ended. The following example is based on [28, Example 2.1].
@@
Consider the rank-two free group $F=\langle a,b\rangle$, thought of as the fundamental group of the wedge of two circles $\Gamma=S^{1}\vee S^{1}$. Consider the element $w=abab^{2}ab^{3}$ of $F$ and realise the corresponding cyclic word by an immersion of a circle $\gamma:S^{1}\to\Gamma$. Let $X$ be the non-positively curved square complex obtained by gluing either end of a cylinder $S^{1}\times[0,1]$ to two copies of $\Gamma$, using $\gamma$ as the attaching map at each end. The links of the two vertices of $X$ are both isomorphic to the Whitehead graph of $w$, which has a cut vertex.
@@
by the Seifert–van Kampen theorem. Applying the automorphism $\phi$ of $F_{2}$ setting $\phi(a)=ab^{-2}$ and $\phi(b)=b$ sends $w\mapsto w^{\prime}=ab^{-1}a^{2}b$. We may realise $w^{\prime}$ by an immersion $\gamma^{\prime}:S^{1}\to\Gamma$ and glue along a cylinder as before to obtain a new square complex $X^{\prime}$ with $\pi_{1}(X^{\prime})\cong\pi_{1}(X)$. However, the links of the two vertices of $X^{\prime}$ are isomorphic to the Whitehead graph of $w^{\prime}$, which consists of two triangles glued along an edge, and in particular is connected with no cut vertices. Therefore, by Theorem E, $\pi(X)\cong\pi_{1}(X^{\prime})$ is one-ended, even though the links of $X$ have cut vertices.
@@
The existence of $X^{\prime}$ is [53, Lemma 4.6] stated for square complexes (see also [31, §3]); it remains to prove that $X^{\prime}$ is non-positively curved. To do this, it suffices to check Gromov’s link condition, that any cycle $\gamma$ in a link of a vertex of $X^{\prime}$ should have length at least $2\pi$ in the angular metric [8, Lemma II.5.6]. Because $\phi$ is a homeomorphism on the complement of the 1-skeleton, the induced maps on links are injective on edges and hence immersions. Thus, $\gamma$ is also an immersed cycle in a link of $X$, so has length at least $2\pi$ by the link condition for $X$. ∎
@@
and each $\pi_{1}(X_{i})$ is non-trivial and freely indecomposable. Furthermore, the unfolding procedure that produces $X^{\prime}$ from $X$ is algorithmic: one simply unfolds an edge of the 1-skeleton whenever the link of a vertex contains a cut vertex. In summary, Theorem E and Proposition 2.29 together provide an effective, geometric method to determine the Grushko decomposition of the fundamental group of a non-positively curved square complex.
@@ 2.6 Shepherd’s theorem and $1$-cuts
A quarter-space of a CAT(0) cube complex is the intersection of two half-spaces. A quarter-space is said to be deep if it is not contained in a bounded neighbourhood of the intersection of its two defining hyperplanes. The next result is [42, Theorem 1.5].
@@
We claim that $U$ is freely essential. Indeed, suppose that some hyperplane $K$ of $U$ had a compact half-space $V$. If $K$ were disjoint from $H$ then $K$ would also be an inessential hyperplane in $\widetilde{X}$, contradicting the hypothesis that $\widetilde{X}$ is essential. If $K$ crossed $H$ then $V$ would be a compact quarter-space in $\widetilde{X}$, contradicting the hypothesis that all quarter-spaces are deep.
@@ 3.1 Periodic $2$-cuts
Let $X$ be compact, non-positively curved cube complex. An abstract hyperplane component over $X$ consists of the following data:
@@
    a non-trivial partition of $\pi_{0}(\mathrm{Wh}_{H}(Z))$.
Two abstract hyperplane components $Z\subseteq H$ and $Z^{\prime}\subseteq H^{\prime}$ are of isomorphic type if there is a deck transformation $\phi\in\pi_{1}(X)$ such that $\phi(Z)=Z^{\prime}$, inducing an isomorphism $\mathrm{Wh}_{H}(Z)\cong\mathrm{Wh}_{H^{\prime}}(Z^{\prime})$ that preserves the chosen partitions of $\pi_{0}$.
@@
Let $X$ be a compact, non-positively curved cube complex. Consider a 2-cut $Y$ in $X$ between hyperplanes $H,H^{\prime}$. Define an oriented abstract hyperplane component corresponding to $H$ as follows:
@@
    the partition on $\pi_{0}(\mathrm{Wh}_{H}(Z))=\pi_{0}(\mathrm{Lk}_{\mathrm{Wh}_{X}(Y)})(H)$ is induced by connectivity in $\mathrm{Wh}_{X}(Y)_{\{H,H^{\prime}\}}$;
@@
Define an oriented abstract hyperplane component corresponding to $H^{\prime}$ in the same way. The 2-cut $Y$ is said to be periodic if the two oriented abstract hyperplane components are of opposite type. In this case, the element $\phi\in\pi_{1}(X)$ sending $H\to H^{\prime}$ is called the associated covering transformation of $Y$.
@@
Choosing a base vertex * in $\widetilde{X}$, the condition that $e(G,C)>1$ means that, for some $R$, the $R$-neighbourhood of the orbit $C*$ separates $\widetilde{X}$ into at least two infinite components. Let $\widetilde{Z}$ be the cubical convex hull of the $R$-neighbourhood of $C*$, which still separates $\widetilde{X}$ into at least two infinite components. By Haglund’s theorem (Theorem 1.7), $C$ acts cocompactly on $\widetilde{Z}$. Because $\widetilde{Z}$ separates $\widetilde{X}$ into two components, $\mathrm{Wh}_{X}(\widetilde{Z})$ is disconnected by Lemma 2.12, as required.
Conversely, suppose that some $C\cong\mathbb{Z}$ acts cocompactly on some convex subcomplex $\widetilde{Z}$ with $\mathrm{Wh}_{X}(\widetilde{Z})$ disconnected, so $\widetilde{Z}$ separates $\widetilde{X}$ into two components by Lemma 2.12. Either some complementary component remains within a bounded distance of $\widetilde{Z}$, contradicting the hypothesis that $X$ is essential, or all complementary components are unbounded away from $\widetilde{Z}$, meaning that $e(G,C)>1$, as required. ∎
@@
which is $\langle\phi\rangle$-invariant and cocompact. By Proposition 2.21, $\mathrm{Wh}_{X}(\widetilde{Z})$ is constructed by iteratively splicing copies of $\mathrm{Wh}_{X}(Y)$ along copies of $H$, meaning that $\mathrm{Wh}_{X}(\widetilde{Z})$ is the infinite union
@@
where $L=\mathrm{Lk}_{\mathrm{Wh}_{X}(Y)}(H)\cong\mathrm{Wh}_{H}(H_{Y})$ by Lemma 2.13. The partitions that are part of the data of an abstract hyperplane component imply that there is a class $c\in\tilde{H}_{0}(L)$ with the property that $c$ and $\phi_{*}(c)$ have the same image in $\tilde{H}_{0}(\mathrm{Wh}_{X}(Y)_{\{H,H^{\prime}\}})$. Applying Mayer–Vietoris and using that homology commutes with direct limits, it follows that $\tilde{H}_{0}(\mathrm{Wh}_{X}(\widetilde{Z}))$ is non-trivial, so $\mathrm{Wh}_{X}(\widetilde{Z})$ is indeed disconnected as required.
Conversely, suppose that there is a convex $\mathbb{Z}$-periodic subcomplex $\widetilde{Z}\subseteq\widetilde{X}$ with $\mathrm{Wh}_{X}(\widetilde{Z})$ disconnected. The compact quotient $Z=\mathbb{Z}\backslash\widetilde{Z}$ has fundamental group $\mathbb{Z}$, and therefore $H^{1}(Z)\cong\mathbb{Z}$. Since the first cohomology is generated by cocycles dual to hyperplanes, some hyperplane $H$ of $Z$ is crossed by any loop representing the generator $\phi$ of $\mathbb{Z}$. Since $\pi_{1}(H)$ is contained in the kernel of the dual map
@@
it follows that $H$ is simply connected, and so lifts to a hyperplane (still denoted by $H$) of $\widetilde{Z}$. By [11, Lemma 2.3], after replacing $\phi$ by a proper power we may assume that $\phi(H)$ is disjoint from $H$.
@@
We claim that $\mathrm{Wh}_{X}(\widetilde{Z}_{+})$ and $\mathrm{Wh}_{X}(\widetilde{Z}_{-})$ are connected. Indeed, $\widetilde{Z}_{+}$ may be written as the ascending union of compact, convex subcomplexes $Y_{n}$ where
@@
and so, since we saw above that $\mathrm{Wh}_{X}(\phi(\widetilde{Z}_{+}))\cong\mathrm{Wh}_{X}(\widetilde{Z}_{+})$ is connected, $Y$ is a 2-cut between $H$ and $H^{\prime}$ by Lemma 2.24. Finally, note that $Y$ is a periodic 2-cut by construction. ∎

### v2

@@ 2.3 Connected sums of simplicial complexes
2.4 Cut-set lemmas in connected sums
@@
Let $C=A\#_{v}B$ for notational convenience. The hypothesis that $\sigma$ is a cut simplex of $C$ means that $C_{\sigma}$ is disconnected. The proof now divides into cases, according to whether $\sigma$ is in $A$ but not in $B$, or $\sigma$ is contained in the canonical copy of the link of $v$.
@@
so Lemma 2.22 implies that $v$ is a cut vertex in $B$, as in item (i), or $A_{\sigma}$ is disconnected, i.e. $\sigma$ is a cut simplex in $A$.
@@ 2.5 Cuts and the generalised Whitehead’s lemma
Consider a finite CAT(0) cube complex $Y$ equipped with an immersion $Y\to X$. If $\mathrm{Wh}_{X}(Y)$ has a finite cut set then $Y$ is said to be a cut. More precisely, if $\mathrm{Wh}_{X}(Y)$ is $(k-1)$-vertex-connected but not $k$-vertex-connected – i.e. if the smallest cut set for $\mathrm{Wh}_{X}(Y)$ has cardinality $k$ – then $Y$ is said to be a $k$-cut. If $\{H_{1},\ldots,H_{k}\}$ is a minimal cut set for a $k$-cut $Y$ then we say that $Y$ is a cut between $H_{1},\ldots,H_{k}$.
@@
Consider a finite CAT(0) cube complex $Y$ equipped with a locally isometric immersion $Y\to X$, and let $Y=Y_{1}\cup_{H}Y_{2}$ for some hyperplane $H$ of $Y$. If $Y$ is a 0-cut then either one of $Y_{1}$ or $Y_{2}$ is a 0-cut, or both $Y_{1}$ and $Y_{2}$ are 1-cuts.
@@
In the base case, $Y$ is a single vertex $y$, so $\mathrm{Wh}_{X}(Y)\cong\mathrm{Lk}_{X}(y)$ by Example 2.6 and there is nothing to prove. If $Y$ has more than one vertex, we may choose a (necessarily separating) hyperplane $H$ of $Y$ and write $Y=Y_{1}\cup_{H}Y_{2}$. By the inductive hypothesis, both $\mathrm{Wh}_{X}(Y_{1})$ and $\mathrm{Wh}_{X}(Y_{2})$ are connected without cut simplices, and therefore so is $\mathrm{Wh}_{X}(Y)$ by Lemma 2.27.
@@
Theorem E provides a convenient certificate that $\pi_{1}(X)$ does not split freely. However, the sufficient condition that it provides is not necessary.
@@
Let $X_{0}$ be the standard 2-torus, seen as a non-positively curved square complex with a single vertex $x_{0}$. Let $X=X_{0}\vee_{x_{0}}I$, where $I$ is a unit interval. The link $\mathrm{Lk}_{X}(x_{0})$ is the disjoint union of a 4-cycle and a point, and in particular is disconnected. However, $\widetilde{X}$ is clearly one-ended.
Examples like 2.28 can be handled by collapsing free edges and free faces. (Free faces manifest themselves as vertices of valence 1 in the link of a vertex.) However, there are also non-positively curved square complexes $X$, with $\widetilde{X}$ one-ended, in which the link of a vertex is connected without valence-1 vertices, but contains a cut vertex.
The following example is based on [36, Example 2.1].
Example 2.29.
Consider the rank-two free group $F=\langle a,b\rangle$, thought of as the fundamental group of the wedge of two circles $\Gamma=S^{1}\vee S^{1}$. Consider the element $w=abab^{2}ab^{3}$ of $F$ and realise the corresponding cyclic word by an immersion of a circle $\gamma:S^{1}\to\Gamma$. Let $X$ be the non-positively curved square complex obtained by gluing either end of a cylinder $S^{1}\times[0,1]$ to two copies of $\Gamma$, using $\gamma$ as the attaching map at each end. The links of the two vertices of $X$ are both isomorphic to the first subdivision of the Whitehead graph of $w$, which is connected, without valence-1 vertices, but has a cut vertex.
@@
by the Seifert–van Kampen theorem. Applying the automorphism $\phi$ of $F_{2}$ setting $\phi(a)=ab^{-2}$ and $\phi(b)=b$ sends $w\mapsto w^{\prime}=ab^{-1}a^{2}b$. We may realise $w^{\prime}$ by an immersion $\gamma^{\prime}:S^{1}\to\Gamma$ and glue along a cylinder as before to obtain a new square complex $X^{\prime}$ with $\pi_{1}(X^{\prime})\cong\pi_{1}(X)$. However, the links of the two vertices of $X^{\prime}$ are isomorphic to the first subdivision of the Whitehead graph of $w^{\prime}$, which consists of two triangles glued along an edge, and in particular is connected with no cut vertices. Therefore, by Theorem E, $\pi(X)\cong\pi_{1}(X^{\prime})$ is one-ended, even though the links of $X$ have cut vertices.
@@
The existence of $X^{\prime}$ is [66, Lemma 4.6] stated for square complexes (see also [39, §3]). The idea of the proof is an algorithmic unfolding procedure: one simply unfolds an edge of the 1-skeleton whenever the link of a vertex contains a cut vertex.
It remains to prove that $X^{\prime}$ is non-positively curved. To do this, it suffices to check Gromov’s link condition, that any cycle $\gamma$ in a link of a vertex of $X^{\prime}$ should have length at least $2\pi$ in the angular metric [10, Lemma II.5.6]. Because $\phi$ is a homeomorphism on the complement of the 1-skeleton, the induced maps on links are injective on edges and hence immersions. Thus, $\gamma$ is also an immersed cycle in a link of $X$, so has length at least $2\pi$ by the link condition for $X$. ∎
@@
and each $\pi_{1}(X_{i})$ is non-trivial and freely indecomposable. Since the proof of Proposition 2.30 is algorithmic, Theorem E and the proposition together provide an effective, geometric method to determine the Grushko decomposition of the fundamental group of a non-positively curved square complex.
@@ 2.6 Shepherd’s theorem and $1$-cuts
A quarter-space of a CAT(0) cube complex is the intersection of two transverse half-spaces. A quarter-space is said to be deep if it is not contained in a bounded neighbourhood of the intersection of its two defining hyperplanes. The next result is [55, Theorem 1.5].
@@
We claim that $U$ is freely essential. Indeed, suppose that some hyperplane $K$ of $U$ had a compact half-space $V$. If $K$ were disjoint from $H$ then $K$ would also be a compact hyperplane in $\widetilde{X}$, contradicting the hypothesis that $\widetilde{X}$ is essential and one-ended. If $K$ crossed $H$ then $V$ would be a compact quarter-space in $\widetilde{X}$, contradicting the hypothesis that all quarter-spaces are deep.
@@ 3.1 Periodic $2$-cuts
Let $X$ be a compact, non-positively curved cube complex. An abstract hyperplane component over $X$ consists of the following data:
@@
    a non-trivial class $c_{Z}$ in the reduced 0-cohomology $\tilde{H}^{0}(\mathrm{Wh}_{H}(Z);\mathbb{Z}/2\mathbb{Z})$.
Two abstract hyperplane components $Z\subseteq H$ and $Z^{\prime}\subseteq H^{\prime}$ are of isomorphic type if there is a deck transformation $\phi\in\pi_{1}(X)$ such that $\phi(Z)=Z^{\prime}$, inducing an isomorphism $\mathrm{Wh}_{H}(Z)\cong\mathrm{Wh}_{H^{\prime}}(Z^{\prime})$ under which $c_{Z^{\prime}}$ pulls back to $c_{Z}$.
@@
Let $X$ be a compact, non-positively curved cube complex without 0- or 1-cuts. Consider a 2-cut $Y$ in $X$ between hyperplanes $H,H^{\prime}$. Fix a non-trivial class $c_{Y}\in\tilde{H}^{0}(\mathrm{Wh}_{X}(Y)_{\{H,H^{\prime}\}};\mathbb{Z}/2\mathbb{Z})$. Define an oriented abstract hyperplane component corresponding to $H$ as follows:
@@
    set $c_{Z}$ to be the pullback of $c_{Y}$ under the inclusion of $\mathrm{Wh}_{H}(Z)\cong\mathrm{Lk}_{\mathrm{Wh}_{X}(Y)}(H)$ into $\mathrm{Wh}_{X}(Y)_{\{H,H^{\prime}\}}$;
@@
Define an oriented abstract hyperplane component $Z^{\prime}$ corresponding to $H^{\prime}$ in the same way. The 2-cut $Y$ is said to be periodic if the two oriented abstract hyperplane components are of opposite type, for some choice of $c_{Y}$. In this case, the element $\phi\in\pi_{1}(X)$ sending $H\to H^{\prime}$ is called the associated covering transformation of $Y$.
Remark 3.3.
Since $X$ has no 1-cuts, $\mathrm{Wh}_{X}(Y)_{\{H^{\prime}\}}$ is connected. The Mayer–Vietoris sequence applied to the decomposition
$\mathrm{Wh}_{X}(Y)_{\{H^{\prime}\}}=\mathrm{Wh}_{X}(Y)_{\{H,H^{\prime}\}}\cup_{\mathrm{Lk}(H)}C(\mathrm{Lk}(H))$
then implies that the natural map
$\tilde{H}^{0}(\mathrm{Wh}_{X}(Y)_{\{H,H^{\prime}\}};\mathbb{Z}/2\mathbb{Z})\to\tilde{H}^{0}(\mathrm{Wh}_{H}(Z);\mathbb{Z}/2\mathbb{Z})$
is injective, so $c_{Z}$ is indeed non-trivial, as required.
@@
Choosing a base vertex * in $\widetilde{X}$, the condition that $e(G,C)>1$ implies that, for some $R$, the $R$-neighbourhood of the orbit $C*$ separates $\widetilde{X}$ into at least two infinite components. Let $\widetilde{Z}$ be the cubical convex hull of the $R$-neighbourhood of $C*$, which still separates $\widetilde{X}$ into at least two infinite components. By Haglund’s theorem (Theorem 1.7), $C$ acts cocompactly on $\widetilde{Z}$. Because $\widetilde{Z}$ separates $\widetilde{X}$ into two components, $\mathrm{Wh}_{X}(\widetilde{Z})$ is disconnected by Lemma 2.12, as required.
Conversely, suppose that some $C\cong\mathbb{Z}$ acts cocompactly on some convex subcomplex $\widetilde{Z}$ with $\mathrm{Wh}_{X}(\widetilde{Z})$ disconnected, so $\widetilde{Z}$ separates $\widetilde{X}$ into two components by Lemma 2.12. Either some complementary component remains within a bounded distance of $\widetilde{Z}$, in which case some bounding hyperplane of $\widetilde{Z}$ is inessential, which is a contradiction, or all complementary components are unbounded away from $\widetilde{Z}$, meaning that $e(G,C)>1$, as required. ∎
@@
which is $\langle\phi\rangle$-invariant and cocompact. By Proposition 2.21, $\mathrm{Wh}_{X}(\widetilde{Z})$ is constructed by iteratively taking connect sums of copies of $\mathrm{Wh}_{X}(Y)$ along copies of $H$, meaning that $\mathrm{Wh}_{X}(\widetilde{Z})$ is the infinite union
@@
where $L=\mathrm{Lk}_{\mathrm{Wh}_{X}(Y)}(H)\cong\mathrm{Wh}_{H}(H_{Y})$ by Lemma 2.13. The Mayer–Vietoris sequence for reduced cohomology (with coefficients in $\mathbb{Z}/2\mathbb{Z}$) asserts that
$0\to\tilde{H}^{0}(\mathrm{Wh}_{X}(\widetilde{Z}))\to\prod_{n\in\mathbb{Z}}\tilde{H}^{0}(\mathrm{Wh}_{X}(Y)_{\{H,H^{\prime}\}})\to\prod_{n\in\mathbb{Z}}\tilde{H}^{0}(L)$
is exact. The right-hand map sends the non-trivial element
$(c_{Y})_{n\in\mathbb{Z}}\in\prod_{n\in\mathbb{Z}}\tilde{H}^{0}(\mathrm{Wh}_{X}(Y)_{\{H,H^{\prime}\}})$
to $(c_{Y})-(c_{Y})=0$, so $\tilde{H}^{0}(\mathrm{Wh}_{X}(\widetilde{Z}))$ is non-trivial, and $\mathrm{Wh}_{X}(\widetilde{Z})$ is indeed disconnected as required.
Conversely, suppose that there is a convex $\mathbb{Z}$-periodic subcomplex $\widetilde{Z}\subseteq\widetilde{X}$ with $\mathrm{Wh}_{X}(\widetilde{Z})$ disconnected. The compact quotient $Z=\mathbb{Z}\backslash\widetilde{Z}$ has fundamental group $\mathbb{Z}$, and therefore $H^{1}(Z;\mathbb{Z})\cong\mathbb{Z}$. Since the first cohomology is generated by cocycles dual to hyperplanes, some hyperplane $H$ of $Z$ is crossed by any loop representing the generator $\phi$ of $\mathbb{Z}$. Since $\pi_{1}(H)$ is contained in the kernel of the dual map
@@
it follows that $H$ is simply connected, and so lifts to a hyperplane (still denoted by $H$) of $\widetilde{Z}$. By [14, Lemma 2.3], after replacing $\phi$ by a proper power we may assume that $\phi(H)$ is disjoint from $H$.
@@
We claim that $\mathrm{Wh}_{X}(\widetilde{Z}_{+})$ and $\mathrm{Wh}_{X}(\widetilde{Z}_{-})$ are connected. Indeed, let $Y$ be the compact component of $\widetilde{Z}$ that results from cutting along $H$ and $\phi(H)$. Then $\widetilde{Z}_{+}$ may be written as the ascending union of compact, convex subcomplexes $Y_{n}$ where
@@
and so, since we saw above that $\mathrm{Wh}_{X}(\phi(\widetilde{Z}_{+}))\cong\mathrm{Wh}_{X}(\widetilde{Z}_{+})$ is connected, $Y$ is a 2-cut between $H$ and $H^{\prime}$ by Lemma 2.24.
Finally, note that $Y$ is a periodic 2-cut. Indeed, setting $H^{\prime}=\phi(H)$, it only remains to construct a suitable $c_{Z}\in\tilde{H}(Z)$ that satisfies Definition 3.2(ii). To do this, as above we consider the exact sequence
$0\to\tilde{H}^{0}(\mathrm{Wh}_{X}(\widetilde{Z}))\to\prod_{n\in\mathbb{Z}}\tilde{H}^{0}(\mathrm{Wh}_{X}(Y)_{\{H,H^{\prime}\}})\to\prod_{n\in\mathbb{Z}}\tilde{H}^{0}(L)$
provided by Mayer–Vietoris, where $L=\mathrm{Wh}_{H}(H_{Y})$. Since $\mathrm{Wh}_{X}(\widetilde{Z})$ is disconnected, there is a non-trivial class $a\in\tilde{H}^{0}(\mathrm{Wh}_{X}(\widetilde{Z}))$. By injectivity, the image $(b_{n})_{n\in\mathbb{Z}}$ is non-zero, so we may set $c_{Y}=b_{n}$ for any $n$ such that $b_{n}\neq 0$. Finally, exactness of the middle term implies that item (ii) of Definition 3.2 is satisfied. ∎